OpenStudy (anonymous):

pants/shirt cost $71.90. shirt cost $20 less than pants, what's the cost of the shirt? what's cost of pants/ my answer: shirt $14.38 pants $57.52

OpenStudy (anonymous):

that cannot be right because 57.52 is not 20 more than 14.38

OpenStudy (anonymous):

put say x = amount of shirt, so \[x+20\] is the price of the pants and then solve \[x+x+20=71.90\]

OpenStudy (anonymous):

sure first perhaps write \[2x+20=71.90\] then subtract 20 from both sides. what do you get?

OpenStudy (anonymous):

$ 29.90 shirt and pants $42.00

OpenStudy (anonymous):

again that cannot be correct, because the pants are $20 more than the shirt and 42 is not 20 more than 29.90

OpenStudy (anonymous):

\[2x+20=71.90\] \[2x=51.90\] \[x=51.90\div 2\] \[x=25.95\]

OpenStudy (anonymous):

shirt is $25.95, pants are $45.95

OpenStudy (anonymous):

thank you, I tried, thanks for not just giving up the answer you made me try to get the answer for myself, then you gave me the right answer

OpenStudy (anonymous):

where is this symbol on the computer: division?